# Targeted synthetic data generation for tabular data via hardness characterization

### v1

@@ Abstract
Synthetic data generation has been proven successful in improving model performance and robustness in the context of scarce or low-quality data. Using the data valuation framework to statistically identify beneficial and detrimental observations, we introduce a novel augmentation pipeline that generates only high-value training points based on hardness characterization. We first demonstrate via benchmarks on real data that Shapley-based data valuation methods perform comparably with learning-based methods in hardness characterisation tasks, while offering significant theoretical and computational advantages. Then, we show that synthetic data generators trained on the hardest points outperform non-targeted data augmentation on simulated data and on a large scale credit default prediction task. In particular, our approach improves the quality of out-of-sample predictions and it is computationally more efficient compared to non-targeted methods.
@@ 1 Introduction
Training complex machine learning models requires large amounts of data, but in real-world applications data may be of poor quality, insufficient in amount, or subject to privacy, safety, and regulatory limitations. Such challenges have sparked an interest in synthetic data generation (SDG), representing the practice of using available data to generate realistic synthetic samples (Lu et al., 2024). In this work, we argue that, when the objective is to use synthetic data to make an existing machine learning model better generalize to unseen data, augmenting only the hardest training points is more effective than augmenting the entire training dataset. In other words, we propose a scalable targeted synthetic data generation framework, focusing on binary classification tasks. The underlying intuition is that, within a dataset, some of the observations are obvious to classify, whereas others play a more significant role in determining the decision boundary of the trained model. Focusing only on these harder examples when generating additional data may allow the model to efficiently learn a more robust decision boundary.
@@
In particular, the method proposed in this work is specifically targeted towards binary classifiers on tabular data, which often require adaptations such as oversampling, undersampling, or synthetic data generation to improve model performance (Chawla et al., 2002; He et al., 2008) under class imbalance. While deep neural networks have proven successful across domains such as images, audio or text, they are still regularly outperformed on tabular data by simpler and more interpretable tree-based architectures (Shwartz-Ziv & Armon, 2021). Therefore, we choose trees as the reference model for binary classification on tabular data, while deep learning models are used for synthetic data generation. In general, we will use $D=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ to denote the training data, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector and $y_{i}\in\{0,1\}$ is a binary label.
@@
Our targeted synthetic data generation approach is divided into two main steps: (i) hardness characterization of training data points, followed by (ii) training of a synthetic data generator only on the hardest data points of the initial training dataset. In order to identify “hard” data points, we propose a novel hardness characterizer specific to tabular data, based on KNN data Shapleys (Jia et al., 2020), capable of achieving performances comparable to state-of-the-art methods in hardness detection benchmarks, while offering key advantages such as being deterministic, model agnostic and computationally efficient. Hardness characterization has already been applied in the literature with rationales similar to this work, but mainly with the objective of pruning the easier examples (Jia et al., 2021; Paul et al., 2023), rather than augmenting the harder ones.
@@
Once the hardest points have been identified, synthetic data generation models are trained only on the most difficult points identified by our hardness characterizer. In this work, we focus on two widely used deep learning architectures for synthetic data generation for tabular data, namely Tabular Variational Autoencoders (TVAE) and Conditional Tabular Generative Adversarial Networks (CTGAN) (both introduced in Xu et al., 2019). To mitigate instabilities encountered during training, we devise a custom early stopping condition which allows to interrupt training when the quality of the synthetically generated samples is maximised. The best synthesizers are then found using a highly efficient routine which combines model selection and hyperparameter tuning. From a comparison of the best models we find that, given a fixed number of points to generate, augmenting only hard data results in a larger performance improvement on unseen data with respect to non-targeted data augmentation, while being computationally cheaper.
@@
The rest of the work is structured as follows: after a recap on the related literature in Section 2, the proposed approach is discussed and empirically verified in Section 3. Next we explore the use of KNN Shapleys as hardness characterizers via a benchmark study in Section 4.1, followed by a discussion on how to utilize them for synthetic data generation on real (Section 4.2) and simulated data (Section 4.3). The results are extensively discussed in Section 5.
@@ 2 Background and related literature
In this work, we aim to bridge the gap between existing methods for hardness characterization and game theoretic approaches to data valuation, with the objective of improving the performance of synthetic data generators. Therefore, our work closely aligns with existing techniques in hardness characterization, data valuation, and synthetic data generation, briefly reviewed in this section.
@@ 2.1 Hardness characterization
As remarked in the introduction, the first aim of this work is the identification of a subset of points that are hard for the model to learn. Existing literature on the topic of hardness characterization is often qualitative and based on different definitions of “hardness”. Following the taxonomy presented in Seedat et al. (2024), three main categories of hardness can be identified: (i) Mislabeling: the true label is replaced with an incorrect one. The label perturbation probability can be uniform across classes or label-dependent; (ii) Out-of-distribution: features are transformed or shifted, leading to observations distributed differently from the main data generating process; (iii) Atypical: observations that are compatible with the main data generating process, but located in tails or rare regions of the feature distributions. These distinctions normally dictate different courses of actions: datapoints that are mislabeled or out of distribution are symptomatic of errors in the data collection process and should be removed from training data, while atypical datapoints are valid points and may in fact need to be augmented to train a robust model (Yuksekgonul et al., 2023).
While many different hardness characterizers have been proposed, most of them are based around keeping track of some statistic of interest for each datapoint during the training process, meaning they can be adapted to any model that is trained iteratively (such as XGBoost or a neural network). Examples include GraNd (Paul et al., 2023), which uses large gradients as an indicator of hardness, and Forget (Toneva et al., 2019), which counts how many times each point is learnt and then forgotten during training, with the simple intuition that hard points are forgotten more often. Based on the quantitative benchmarks reported in Seedat et al. (2024), Data-IQ (Seedat et al., 2022) emerges as one of the best performers on tabular data, and it is chosen as benchmark hardness characterizer in this work because of its interpretability and compatibility with XGBoost.
Data-IQ
Data-IQ (Seedat et al., 2022) partitions the training data into “easy” and “hard” points via the concepts of epistemic and aleatoric uncertainty, combined with a model confidence score calculated from $\mathcal{P}(\bm{x},\bm{\theta})$, representing the probability of assigning a correct label to an observation, given input features $\bm{x}$ and model parameters $\bm{\theta}$. The epistemic uncertainty is caused by the variability in the estimated model parameters, whereas the aleatoric uncertainty captures the inherent data uncertainty. Given parameter estimates $\{\bm{\theta}_{e}\}_{e=1}^{E}$ obtained during training, a Data-IQ confidence score is obtained as $\bar{\mathcal{P}}(\bm{x})=E^{-1}\sum_{e=1}^{E}\mathcal{P}(\bm{x},\bm{\theta}_{e})$, whereas the aleatoric uncertainty is given by $v_{\mathrm{al}}(\bm{x})=E^{-1}\sum_{e=1}^{E}\mathcal{P}(\bm{x},\bm{\theta}_{e})[1-\mathcal{P}(\bm{x},\bm{\theta}_{e})]$. Training data can then be partitioned into (i) easy points, corresponding to high $\bar{\mathcal{P}}(\bm{x})$ and low $v_{\mathrm{al}}(\bm{x})$, (ii) hard points, corresponding to low $\bar{\mathcal{P}}(\bm{x})$ and low $v_{\mathrm{al}}(\bm{x})$, and (iii) ambiguous points, with high $v_{\mathrm{al}}(\bm{x})$.
@@ 2.2 Data valuation
Data valuation (see Ghorbani & Zou, 2019) is the task of equitably quantifying the contribution of each data point to a model, via a data value $\phi_{i}(D,\mathcal{A},V)$, where $D$ is the training dataset, $\mathcal{A}$ is a model and $V$ is a performance score, where $V(S)$ is the performance score of $\mathcal{A}$ when trained on $S\subseteq D$. Ghorbani & Zou (2019) identifies that a suitable data valuator is the data Shapley of the $i$-th training point, which takes the following form:
@@
where $C$ is an arbitrary constant. The valuator $\phi_{i}$ can be interpreted as weighted sum of all the possible “marginal contributions” of the datapoint. Exact evaluation of the data Shapleys is prohibitively expensive because of the need to retrain the model $2^{n-1}$ times (once per subset) for each observation in the training data. It is possible to approximate the data Shapleys by considering only a small number of subsets $S\subseteq D\setminus\{i\}$ in equation 1, obtained via Monte-Carlo sampling of random permutations. Unfortunately, even state-of-the-art approximations based on this idea, such as TMC-Shapley (Ghorbani & Zou, 2019) or Beta-Shapley (Kwon & Zou, 2022) are computationally prohibitive for datasets with $n\gg 1000$, making them difficult to use in practice. More generally, extensive work has been developed in the literature to make the computation of data Shapleys more efficient, such as Jia et al. (2019), Wang et al. (2023) and Wang et al. (2024).
In particular, Jia et al. (2020) derived a recursive algorithm to compute the exact data Shapleys for KNN classifiers in $\mathcal{O}\{n\log(n)\,n_{\mathrm{test}}\}$ complexity. For training data $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ and test data $\{(\bm{x}_{\mathrm{test},i},y_{\mathrm{test},i})\}_{i=1}^{n_{\mathrm{test}}}$, let $(\alpha_{j,1},\ldots,\alpha_{j,n})$ be the indices of training data points, sorted in increasing order according to their Euclidean distance from the $j$-th test data point, for $j=1,\dots,n_{\mathrm{test}}$. The KNN Shapley for the $i$-th training data point is then obtained as the average $s_{i}=n_{\mathrm{test}}^{-1}\sum_{j=1}^{n_{\mathrm{test}}}s_{j,i}$, where $s_{j,i}$ is calculated recursively as follows:
$s_{j,\alpha_{j,n}}=\frac{1}{n}\mathds{1}_{y_{\mathrm{test},j}}\{y_{\alpha_{j,n}}\},$ $s_{j,\alpha_{j,i}}=s_{j,\alpha_{j,i+1}}+\frac{\mathds{1}_{y_{\mathrm{test},j}}\{y_{\alpha_{j,i}}\}-\mathds{1}_{y_{\mathrm{test},j}}\{y_{\alpha_{j,i+1}}\}}{iK\cdot\min\{K,i\}^{-1}},$ (2)
where $j=1,\dots,n_{\mathrm{test}}$, $i=1,\dots,n-1$, and $\mathds{1}_{\cdot}\{\cdot\}$ is the indicator function. Jia et al. (2021) found KNN Shapleys to be a valid alternative to data Shapleys in tasks such as data summarization or noisy labels detection, only when based on deep features extracted from image data by pre-trained architectures. A similar approach is unfortunately not possible on tabular data. In this work we argue that KNN Shapleys computed using a test dataset large enough to cover the entire data distribution, can be used for hardness characterization on tabular data. Specifically, we argue that the lowest KNN Shapleys identify the hardest training points.
@@ 2.3 Synthetic data generation
Synthetic data generators (SGDs) for tabular data have been actively researched in recent years (see for example Fonseca & Bacao, 2023). Here, we discuss three of the most popular SDGs used in practice: SMOTE, CTGAN, and TVAE.
SMOTE
In Synthetic Minority Over-sampling Technique (SMOTE; Chawla et al., 2002), observations are picked at random among the $K$ nearest neighbours of the same class for each datapoint, and new data is generated by random sampling along the segment connecting the point to the chosen neighbour. The number of neighbours is determined based on the over-sampling ratio.
CTGAN
Tabular GAN (TGAN; Xu & Veeramachaneni, 2018) was the first attempt at using Generative Adversarial Networks (GAN; Goodfellow et al., 2014) to generate tabular data, which was then extended to Conditional TGANs (CTGANs; Xu et al., 2019). A GAN consists of two models: a generator that takes as input random noise and tries to output a point following the data distribution, and a discriminator that given observations (either from the original dataset or synthesized by the generator) outputs the probability of them being fake. Both the generator and discriminator are typically neural networks. In CTGANs, the possible non-Gaussianity and multimodality in the data is handled by fitting a variational Gaussian mixture model to each normalized numeric feature, whereas categorical features are represented as one-hot encoded vectors. Both the generator and the discriminator networks within the underlying GAN are multilayer perceptrons (MLPs). The name of the CTGAN architecture derives from the form of its conditional generator, which allows to generate data conditional on specific values of the discrete features, for better handling imbalanced datasets.
TVAE
A Tabular VAE (TVAE; Xu et al., 2019) is an adaptation for tabular data of a Variational Auto-Encoder (VAE; Kingma & Welling, 2013), using similar preprocessing to CTGAN. A VAE consists in an encoder that learns to represent high-dimensional data into a low-dimensional latent space, and a decoder that reconstructs the compressed representation into the original domain. Both the encoder and decoder networks are normally chosen to be MLPs (Kingma & Welling, 2013).
@@ 3 KNN Shapleys as hardness characterizers for SDGs
In this work, we propose to utilize KNN Shapleys as hardness characterizers, and subsequently generate synthetic data by only augmenting the hardest points. Our proposed pipeline is the following:
@@
    KNN Shapley calculation – For a model $\mathcal{A}$ fitted on training data $D=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, calculate the KNN Shapleys $s_{i},\ i=1,\dots,n$, based on a test set $\{(\bm{x}_{\mathrm{test},i},y_{\mathrm{test},i})\}_{i=1}^{n_{\mathrm{test}}}$.
@@
    Ranking by hardness – Sort the KNN Shapleys in increasing order, with points with lower values representing harder examples.
@@
    Targeted augmentation – Given a synthetic data generator $\mathcal{G}$, perform data augmentation only on the hardest points, based on the ranking of the KNN Shapleys.
By generating synthetic data for difficult examples only, we aim to improve the performance of the model $\mathcal{A}$ specifically on the most challenging parts of the data distribution. The proposed procedure is summarized visually in Figure 1. In the next section, we provide a mathematical intuition on the interpretation of KNN Shapley as hardness charaterizers.
@@
3.1 KNN Shapleys for hardness characterization: a toy example
Consider a mixture of two univariate normals with unit variance, centred at $-1$ and $1$, such that $p(x\mid y=0)=\mathcal{N}(x;-1,1)$ and $p(x\mid y=1)=\mathcal{N}(x;1,1)$. Datapoints are drawn from each distribution with probability $1/2$. Additionally, consider two points $(x_{1},y_{1})=(-1,0)$ and $(x_{2},y_{2})=(1,1)$, representative of each component of the mixture distribution. Consider a third point $(x_{\mathrm{train}},y_{\mathrm{train}})$, with $y_{\mathrm{train}}=0$ without loss of generality. Hence, the training set takes the form $D=\{(-1,0),(1,1),(x_{\mathrm{train}},0)\}$. As $x_{\mathrm{train}}$ increases, it becomes harder to classify correctly under the true underlying data distribution, since the ratio $p(x_{\mathrm{train}}\mid y=0)\ /\ p(x_{\mathrm{train}}\mid y=1)$ decreases, consequently increasing the probability of misclassification. Note that any value of $x_{\mathrm{train}}\in\mathbb{R}$, combined with $x_{1}$ and $x_{2}$, partitions the real line into four different regions.
Any test data point $(x_{\mathrm{test}},y_{\mathrm{test}})\in\mathbb{R}\times\{0,1\}$ induces a ranking of the training data based on the distances $|x_{\mathrm{test}}-x|,\ x\in\{x_{1},x_{2},x_{\mathrm{train}}\}$. For $x_{\mathrm{train}}=0$, the four possible regions with different rankings of the three training points are plotted in Figure 2(b). The $1$NN Shapley value $s_{\mathrm{train}}$ for the point $(x_{\mathrm{train}},y_{\mathrm{train}})=(0,0)$ can be calculated explicitly for all the possible values of $(x_{\mathrm{test}},y_{\mathrm{test}})$. The results are reported in Table 2(b). Since the data distribution is known, the expected value $\mathbb{E}(s_{\mathrm{train}})$ of the $1$NN Shapley for $(x_{\mathrm{train}},y_{\mathrm{train}})=(0,0)$ can be explicitly calculated:
$\mathbb{E}(s_{\mathrm{train}})$ $=\int_{-\infty}^{+\infty}s_{\mathrm{train}}\,p(x_{\mathrm{test}})\ \mathrm{d}x_{\mathrm{test}}=\frac{1}{2}\sum_{y\in\{0,1\}}\int_{-\infty}^{+\infty}s_{\mathrm{train}}\,p(x_{\mathrm{test}}\,|\,y)\ \mathrm{d}x_{\mathrm{test}}\approx 0.209.$ (3)
It must be remarked that the value of $s_{\mathrm{train}}$ depends on $(x_{\mathrm{test}},y_{\mathrm{test}})$, as detailed in Table 2(b) and equation 1. The same procedure as equation 3 can be repeated for any $x_{\mathrm{train}}\in\mathbb{R}$, with results displayed in Figure 2(c). As expected, the $1$NN Shapley decreases in the direction of increasing hardness, suggesting an association between low KNN Shapleys and hard regions of feature space.
(a) Four regions for $x_{\mathrm{test}}$ and ranking of $\{x_{1},x_{\mathrm{train}},x_{2}\}$, for $x_{\mathrm{train}}=0$.

### v2

@@ Abstract
Data augmentation via synthetic data generation has been shown to be effective in improving model performance and robustness in the context of scarce or low-quality data. Using the data valuation framework to statistically identify beneficial and detrimental observations, we introduce a simple augmentation pipeline that generates only high-value training points based on hardness characterization, in a computationally efficient manner. We first empirically demonstrate via benchmarks on real data that Shapley-based data valuation methods perform comparably with learning-based methods in hardness characterization tasks, while offering significant computational advantages. Then, we show that synthetic data generators trained on the hardest points outperform non-targeted data augmentation on a number of tabular datasets. Our approach improves the quality of out-of-sample predictions and it is computationally more efficient compared to non-targeted methods.
@@ 1 Introduction
Training complex machine learning models requires large amounts of data, but in real-world applications, data may be of poor quality, insufficient in amount, or subject to privacy, safety, and regulatory limitations. Such challenges have sparked an interest in synthetic data generators (SDGs), which use available training data to generate realistic synthetic samples [Lu et al., 2024]. In this work, we argue that when the objective is to use synthetic data to make an existing machine learning model better generalize to unseen data, augmenting only the hardest training points is more effective and has significant computational advantages compared to augmenting the entire training dataset. In other words, we propose a novel scalable targeted synthetic data generation framework, focusing on binary classification tasks. The underlying intuition is that within a dataset, some of the observations are obvious to classify, whereas others play a more significant role in determining the decision boundary of the trained model. Focusing only on these harder examples when generating additional data may allow the model to efficiently learn a more robust decision boundary with the additional advantage of reducing computational costs associated with the training of a synthetic data generator.
@@
In particular, the method proposed in this work is specifically targeted toward binary classifiers on tabular data, which often require adaptations such as oversampling, undersampling, or synthetic data generation to improve model performance under class imbalance [Chawla et al., 2002, He et al., 2008]. Although deep neural networks have proven successful across domains such as images, audio, or text, they are still regularly outperformed on tabular data by simpler and more interpretable tree-based architectures [Shwartz-Ziv and Armon, 2021]. Therefore, we choose trees as the reference model for binary classification on tabular data, while deep learning models are used for synthetic data generation. In general, we will use $D=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ to denote the training data, where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a $d$-dimensional feature vector and $y_{i}\in\{0,1\}$ is a binary label.
@@
Our targeted synthetic data generation approach is divided into two main steps: (i) hardness characterization of training data points, followed by (ii) training of a synthetic data generator only on the hardest data points of the initial training dataset. In order to identify “hard” data points, we propose a novel hardness characterizer specific to tabular data, based on KNN data Shapleys [Jia et al., 2020]. We show that KNN data Shapleys are capable of achieving performances comparable to state-of-the-art methods in hardness detection benchmarks, while offering key advantages such as being deterministic, model agnostic and, most importantly, computationally efficient when compared to existing alternatives in the literature. The computational advantages of the proposed data augmentation pipeline mainly stem from two aspects: (i) the identification of “hard” data points via KNN Shapleys, a computationally efficient method against other hardness characterizers; and (ii) the reduction of the number of data points used for training of synthetic data generators, focusing only on such “hard” data points. Even though hardness characterization has already been applied in the literature with rationales similar to this work, the focus was mostly on pruning the easier examples [Jia et al., 2021, Paul et al., 2023], rather than augmenting the harder ones.
@@
Once the hardest points have been identified, synthetic data generation models are trained only on the most difficult points identified by our hardness characterizer. In this work, we focus particularly on two widely used deep architectures for synthetic data generation for tabular data, namely Tabular Variational Autoencoders (TVAE) and Conditional Tabular Generative Adversarial Networks (CTGAN) [both introduced in Xu et al., 2019], but we stress that our pipeline can be applied to any synthetic data generator. When comparing the best models on various different datasets given a fixed number of points to generate, we find that augmenting only hard data points results in a larger performance improvement on unseen data with respect to non-targeted data augmentation, while being computationally cheaper.
@@
The rest of the work is structured as follows: after summarizing the related literature in Section 2, the proposed approach is discussed and empirically verified in Section 3. Next, we explore the use of KNN Shapleys as hardness characterizers via a benchmark study in Section 4.1, followed by an outline on how to utilize them for synthetic data generation on real (Sections 4.3 and 4.4) and simulated data (Section 4.2). The results are discussed in Section 5.
@@ 2 Background and related literature
In this work, we aim to bridge the gap between existing methods for hardness characterization and game theoretic approaches to data valuation, with the objective of improving the performance of synthetic data generators. Therefore, our work aligns with existing techniques in hardness characterization, data valuation, and synthetic data generation.
@@ 2.1 Hardness characterization
As described in the introduction, the first aim of this work is the identification of a subset of points that are hard for the model to classify. Existing literature on hardness characterization is often qualitative and based on different definitions of “hardness”. Following the taxonomy presented in Seedat et al. [2024], three main categories of hardness can be identified: (i) Mislabeling: the true label is replaced with an incorrect one. The label perturbation probability can be uniform across classes or label-dependent; (ii) Out-of-distribution: features are transformed or shifted, leading to observations distributed differently from the main data generating process; (iii) Atypical: observations that are compatible with the main data generating process, but located in tails or rare regions of the feature distributions. These distinctions normally dictate different courses of actions: datapoints that are mislabeled or out of distribution are symptomatic of errors in the data collection process and should be removed from training data, whereas atypical data may need to be augmented to train a robust model [Yuksekgonul et al., 2023].
While many different hardness characterizers have been proposed, most of them are based around keeping track of some statistic of interest for each datapoint during the training process, meaning they can be adapted to any model that is trained iteratively (such as XGBoost or a neural network). Examples include GraNd [Paul et al., 2023], which uses large gradients as an indicator of hardness, and Forget [Toneva et al., 2019a], which counts how many times each point is learned and then forgotten during training, with the simple intuition that hard points are forgotten more often. Based on the quantitative benchmarks reported in Seedat et al. [2024], Data-IQ [Seedat et al., 2022] and CleanLab [Northcutt et al., 2021] are often the best performers on tabular data, and they can be considered as benchmark hardness characterizers for tabular data because of their interpretability and compatibility with XGBoost.
@@ 2.2 Data valuation
Data valuation [see Ghorbani and Zou, 2019] is the task of equitably quantifying the contribution of each data point to a model, via a data value $\phi_{i}(D,\mathcal{A},V)$, where $D$ is the training dataset, $\mathcal{A}$ is a model and $V$ is a performance score, where $V(S)$ is the performance score of $\mathcal{A}$ when trained on $S\subseteq D$. Jiang et al. [2023] propose a framework for benchmarking data valuation algorithms, concluding that no method uniformly outperforms others in every task. Therefore, in this work, we focus on a computationally efficient data valuation mechanism, which we propose to use for hardness characterization: KNN Shapley [Jia et al., 2020].
KNN Shapleys correspond to an instance of the data Shapley valuator Ghorbani and Zou [2019]. For the $i$-th training point in $D$, the data Shapley $\phi_{i}$ takes the following form:
@@
where $C$ is an arbitrary constant. The valuator $\phi_{i}$ can be interpreted as weighted sum of all the possible “marginal contributions” of the datapoint. Exact evaluation of the data Shapleys is prohibitively expensive because of the need to retrain the model $2^{n-1}$ times (once per subset) for each observation in the training data. It is possible to approximate the data Shapleys by considering only a small number of subsets $S\subseteq D\setminus\{i\}$ in (1), obtained via Monte-Carlo sampling of random permutations. Unfortunately, even state-of-the-art approximations based on this idea, such as TMC-Shapley [Ghorbani and Zou, 2019] or Beta-Shapley [Kwon and Zou, 2022] are computationally prohibitive for datasets with $n\gg 1000$, making them difficult to use in practice.
An efficient method to calculate exact data Shapleys for KNN classifiers in $\mathcal{O}\{n\log(n)\,n_{\mathrm{test}}\}$ complexity has been derived in Jia et al. [2020]. For training data $\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ and test data $\{(\bm{x}_{\mathrm{test},i},y_{\mathrm{test},i})\}_{i=1}^{n_{\mathrm{test}}}$, let $(\alpha_{j,1},\ldots,\alpha_{j,n})$ be the indices of training data points, sorted in increasing order according to their Euclidean distance from the $j$-th test data point, for $j=1,\dots,n_{\mathrm{test}}$. The KNN Shapley for the $i$-th training data point is defined as $s_{i}=n_{\mathrm{test}}^{-1}\sum_{j=1}^{n_{\mathrm{test}}}s_{j,i}$, where $s_{j,i}$ is calculated recursively as follows:
$s_{j,\alpha_{j,n}}=\frac{1}{n}\mathds{1}_{y_{\mathrm{test},j}}\{y_{\alpha_{j,n}}\},$ (2)
$s_{j,\alpha_{j,i}}=s_{j,\alpha_{j,i+1}}+\frac{\mathds{1}_{y_{\mathrm{test},j}}\{y_{\alpha_{j,i}}\}-\mathds{1}_{y_{\mathrm{test},j}}\{y_{\alpha_{j,i+1}}\}}{iK\cdot\min\{K,i\}^{-1}},$ (3)
where $j=1,\dots,n_{\mathrm{test}}$, $i=1,\dots,n-1$, and $\mathds{1}_{\cdot}\{\cdot\}$ is the indicator function. Jia et al. [2021] found KNN Shapleys to be a valid alternative to data Shapleys in tasks such as data summarization or noisy labels detection, only when based on deep features extracted from image data by pre-trained architectures. A similar approach is unfortunately not possible on tabular data. In this work we argue that KNN Shapleys can be used for hardness characterization on tabular data. Specifically, we argue that the lowest KNN Shapleys identify the hardest training points.
@@ 2.3 Synthetic data generation
Synthetic data generators (SGDs) for tabular data have been actively researched in recent years [see for example Fonseca and Bacao, 2023]. One of the most popular SDGs used in practice is Synthetic Minority Over-sampling Technique [SMOTE; Chawla et al., 2002], which generates new data by random sampling along the segment connecting each data point to a randomly sampled neighbor. More recently, Conditional TGANs (CTGAN), and Tabular VAE (TVAE) were proposed in Xu et al. [2019]. CTGAN consists in a Generative Adversarial Networks [GAN; Goodfellow et al., 2014] with a conditional generator for tabular data which allows to generate data points conditional on specific values of the discrete features, for better handling imbalanced datasets, whereas TVAE is an adaptation for tabular data of a Variational Auto-Encoder [VAE; Kingma and Welling, 2013]. Diffusion and score-based methods have also been recently proposed in the literature [see, for example Kotelnikov et al., 2023, Kim et al., 2023, Zhang et al., 2024], as well as adversarial random forests [Watson et al., 2023].
The objective of this work is not to compare different SDGs, but rather to propose KNN Shapleys as a computationally efficient hardness characterizer, with the objective to extract “hard” data points to use in the training of any synthetic data generator, making the procedure computationally more efficient without performance deterioration. Therefore, for the experiments in this work, we focus on two of the most popular and best performing SDGs in the literature (TVAE and CTGAN). However, we do emphasize that our procedure is applicable to any SDG.
@@ 3 KNN Shapleys as hardness characterizers for SDGs
In this work, we propose a simple pipeline: first, training data are ranked via a hardness characterizer. In particular, we propose to utilize a computationally efficient data valuator as hardness characterizer: KNN Shapleys, which can be calculated efficiently via (2). Second, we train a synthetic data generator using only the hardest points, which provides a significant computational speedup over training with the entire training set. Our proposed pipeline is the following:
@@
    KNN Shapley calculation – For a model $\mathcal{A}$ fitted on training data $D=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$, calculate the KNN Shapleys $s_{i},\ i=1,\dots,n$, based on a test set $\{(\bm{x}_{\mathrm{test},i},y_{\mathrm{test},i})\}_{i=1}^{n_{\mathrm{test}}}$, using the recursion in (2).
@@
    Ranking by hardness – Sort the KNN Shapley values in increasing order, with points with lower values representing harder examples.
@@
    Targeted augmentation – Given a synthetic data generator $\mathcal{G}$, perform data augmentation only on the $\tau\cdot 100\%$ hardest points, based on the ranking of the KNN Shapleys, for a threshold $\tau\in(0,1]$.
By generating synthetic data for difficult examples only, we aim to improve the performance of the model $\mathcal{A}$ specifically on the most challenging parts of the data distribution. The proposed procedure is summarized visually in Figure 1. This data processing approach could be related to the ideas of dataset distillation [Wang et al., 2020] and dataset condensation [Zhao et al., 2021], commonly used in computer vision, aimed at finding compact representations of a large dataset via a small set of informative samples, for the purposes of data-efficient learning. The proposed approach could also be related to the data profiling step in Hansen et al. [2023], which is used to divide training data points in different subsets (easy, ambiguous and hard) via data valuation methods such as CleanLab [Northcutt et al., 2021] or Data-IQ [Seedat et al., 2022]. Separate SDGs are fitted on each segment and the resulting synthetic data are combined, with the purpose of obtaining the best fully synthetic training dataset in terms of statistical fidelity. Therefore, the objective of Hansen et al. [2023] is not data augmentation, but rather to generate a fully synthetic dataset which performs similarly to the original training data, to overcome data privacy issues. On the other hand, we exclusively focus on data augmentation, with the objective of achieving the largest possible performance boost at a small computational cost: we propose to augment the hardest data points provided by the ranking of KNN Shapleys, combining them with the original training dataset.
@@
3.1 KNN Shapleys for hardness characterization: an example
In this section, we provide a mathematical intuition on the interpretation of KNN Shapleys as hardness charaterizers. Consider a mixture of two univariate normals with unit variance, centred at $-1$ and $1$, such that $p(x\mid y=0)=\mathcal{N}(x;-1,1)$ and $p(x\mid y=1)=\mathcal{N}(x;1,1)$. Datapoints are drawn from each distribution with probability $1/2$. Additionally, consider two points $(x_{1},y_{1})=(-1,0)$ and $(x_{2},y_{2})=(1,1)$, representative of each component of the mixture distribution. Consider a third point $(x_{\mathrm{train}},y_{\mathrm{train}})$, with $y_{\mathrm{train}}=0$ without loss of generality. Hence, the training set takes the form $D=\{(-1,0),(1,1),(x_{\mathrm{train}},0)\}$. As $x_{\mathrm{train}}$ increases, it is harder to classify correctly under the true data distribution, since the ratio $p(x_{\mathrm{train}}\mid y=0)\ /\ p(x_{\mathrm{train}}\mid y=1)$ decreases monotonically (converging to zero in the limit), consequently increasing the probability of misclassification.
Any test data point $(x_{\mathrm{test}},y_{\mathrm{test}})\in\mathbb{R}\times\{0,1\}$ induces a ranking of the training data based on the distances $|x_{\mathrm{test}}-x|,\ x\in\{x_{1},x_{2},x_{\mathrm{train}}\}$. Assuming $x_{\mathrm{train}}=0$, the $1$NN Shapley value $s_{\mathrm{train}}$ for the point $(x_{\mathrm{train}},y_{\mathrm{train}})=(0,0)$ can be calculated explicitly for all the possible values of $(x_{\mathrm{test}},y_{\mathrm{test}})$. The results are reported in Table 2(a). Since the data distribution is known, the expected value $\mathbb{E}(s_{\mathrm{train}})$ of the $1$NN Shapley for $(x_{\mathrm{train}},y_{\mathrm{train}})=(0,0)$ can also be explicitly calculated:
$\mathbb{E}(s_{\mathrm{train}})=\int_{-\infty}^{+\infty}s_{\mathrm{train}}\,p(x_{\mathrm{test}})\ \mathrm{d}x_{\mathrm{test}}\\ =\frac{1}{2}\sum_{y\in\{0,1\}}\int_{-\infty}^{+\infty}s_{\mathrm{train}}\,p(x_{\mathrm{test}}\,|\,y)\ \mathrm{d}x_{\mathrm{test}}\approx 0.209.$ (4)
The value of $s_{\mathrm{train}}$ depends on $(x_{\mathrm{test}},y_{\mathrm{test}})$, as detailed in Table 2(a) and (1). The same procedure as (4) can be repeated for any $x_{\mathrm{train}}\in\mathbb{R}$, with results displayed in Figure 2(b). As expected, the $1$NN Shapley decreases in the direction of increasing hardness, suggesting an association between low KNN Shapleys and hard regions of feature space.